OpenStudy (anonymous):

If the sequence {a^2} converges then the sequence {a} converges true or false and why.

OpenStudy (anonymous):

From logic you may also know: \[\Large p \Rightarrow q \Longleftrightarrow \neg q \Rightarrow \neg p \] So if {a} diverges, then a^2 diverges.

OpenStudy (anonymous):

Sometimes showing that the contrapositive holds true is easier than showing that the implication holds.

OpenStudy (experimentx):

\[ \lim_{n \to \infty }|a_n^2 - L|=\lim_{n \to \infty }|(|a_n| - \sqrt L)(|a_n| + \sqrt L)| = \lim_{n\to \infty }|(|a_n| + \sqrt L)| |(|a_n| - \sqrt L)| \\ \implies \lim_{n\to \infty } (|a_n| - \sqrt L)| < \epsilon/ \sqrt L \]

OpenStudy (experimentx):

since the sequence converges absolutely, it must converge ..

OpenStudy (anonymous):

Just for the sake of completeness, here is the contrapositive. \[\Large \neg q \Rightarrow \neg p \] So we assume that \(a_n\) diverges, as given by the implication of the contrapositive, this means that: \[\Large \lim_{n \to \infty} a_n=\infty\] now we can square both sides to optain: \[\Large \lim_{n \to \infty} a_n^2=\infty^2=\infty \] so we have shown that \(a_n^2\) diverges from our assumption that \(a_n\) diverges, therefore we can state: \[\Large \neg q \Rightarrow \neg p \Longleftrightarrow p \Rightarrow q \] which shows that if \(a_n^2\) converges, so does \(a_n\)

OpenStudy (experimentx):

seems that this would work too ...